If |Z1|=|z2|=|z3|=|z4| and Z1+z2+z3+z4=0, then z1,z2,z3,z4 represent the vertices of_____

Grade:12

1 Answers

Chunduru Vaishnav
6 years ago
Since magnitude​s are equal 
Z1= x+it
Z1= x-iy
Z3= -x-iy
Z4= -x+it
And there sim is zero
And magnitude of all of them is (x2 +y2)1/2
These always form a rectangle